OpenStudy (anonymous):

Let f(x) = 13 - 11x - 5x^2 and g(x) = 11 + 25x - 3x^2 + 4x^3. Then f(x) + g(x) equals

OpenStudy (jmark):

f(x) +g(x) = 13 - 11x – 5x^2 + 11 + 25x - 3x^2 + 4x^3 add the like term. = 24-14x+x^2 ICSE question papers for class 9 - http://icse.edurite.com/icse-question-papers/icse-question-papers-for-class-9.html